From your house you jog 10m in the direction of 25 degree north of east, then adter you changed your direction by 6m, 15 degree north of east. Graph and calculate the magnitude and angle of the resultant vector using analytical method.
1
Expert's answer
2021-08-10T08:00:42-0400

10 m at 25° N of E:


"d_{1N}=10\sin25°=4.23\text{ m},\\
d_{1E}=10\cos25°=9.06\text{ m}."

6 m at 15° N of E:


"d_{2N}=6\sin15°=1.55\text{ m},\\
d_{2E}=6\cos15°=5.80\text{ m}."

The resultant vector components:


"d_N=d_{1N}+d_{2N}=5.78\text{ m}.\\
d_E=d_{1E}+d_{2E}=14.9\text{ m}.\\
d=\sqrt{d_N^2+d_E^2}=15.9\text { m}.\\\space\\
\theta=\arctan{\frac{d_N}{d_E}}
=21.2°."